OpenStudy (anonymous):

How do you write 4 log m-log n into a single logarithm?

OpenStudy (mathmale):

Would you mind typing out the three rules of logs that apply to this kind of problem? Hint: log m + log n = log mn. (Can you put that into words?)

OpenStudy (anonymous):

remember these rules \[\log x - \log y = \log(\frac{ x }{ y })\] \[x \log y = \log y^{x}\]

OpenStudy (mathmale):

Chris: In your shoes I'd take careful notes on this material for later reference. Which of the 3 rules applies to the expression 4 log m? Please explain your reasoning and rewrite 4 log m as appropriate.